# KEK-TH-2773 Quantum vacuum energy and geometry of extra dimension

### v1

@@ 1 Introduction
The paper is organized as follows. In the next section, we provide a brief review of the calculation of the vacuum energy density and the pressure in 4D effective theory, and derive the conditions required to cancel the $\Lambda_{\rm cut}$-dependent part. In Sec. 3, we calculate the VEV of the 5D energy-momentum tensor $\left\langle T_{MN}\right\rangle$ by using the 5D propagators. In Sec. 4, we discuss the dependence of $\left\langle T_{MN}\right\rangle$ on the cutoff energy scale $\Lambda_{\rm cut}$, and its cancellation among bosons and fermions. Sec. 5 is devoted to the summary and discussions. We have collected our notations in App. A, the definitions of the basis functions for solutions to the mode equations in App. B, and the derivation of the 5D propagators in App. C. In App. D, we derive approximate expressions of various quantities for large momentum.
@@ 2 Vacuum energy density in 4D effective theory
When we consider the cosmological evolution of the spacetime with a large extra dimension, such as the dark dimension scenario, we need to solve the 5D field equations because the temperature of the universe can be higher than the KK scale, which is the cutoff scale of the 4D effective theory.
In the following, we consider a 5D setup whose metric is given by
$ds^{2}$ $=g_{MN}dx^{M}dx^{N}=-n^{2}(t,y)dt^{2}+a^{2}(t,y)\sum_{i=1}^{3}(dx^{i})^{2}+b^{2}(t,y)dy^{2},$ (3.1)
where $t\equiv x^{0}$, $y\equiv x^{4}$, and $0\leq y\leq L\equiv\pi R$. Since we are interested in a situation that 4D Friedmann-Lemaitre-Robertson-Walker (FLRW) universe with zero 3D curvature is realized at low energies, we focus on a case that
$a(t,y)$ $=a_{\rm sc}(t)n(t,y),$ (3.2)
where $a_{\rm sc}(t)$ is the scale factor.
The evolution of the spacetime is determined by the Einstein equation, [^6] We adopt the semiclassical approach, in which the spacetime is treated classically, while the quantum effects of matter fields are incorporated via $\left\langle T_{MN}\right\rangle$.
${\cal R}_{MN}-\frac{1}{2}g_{MN}{\cal R}$ $=\kappa_{5}\left\langle T_{MN}\right\rangle,$ (3.3)
where ${\cal R}_{MN}$ and ${\cal R}\equiv g^{MN}{\cal R}_{MN}$ are the 5D Ricci tensor and scalar, respectively, $\kappa_{5}$ is the 5D gravitational coupling, and $\left\langle T_{MN}\right\rangle$ is the expectation value of the energy-momentum tensor $T_{MN}$ for quantum fields. In the following, we calculate $\left\langle T_{MN}\right\rangle$ for scalar and spinor fields.
When we discuss the time evolution of the universe, we need $\left\langle T_{MN}\right\rangle$ at each time. Since the background space expands slowly enough compared to the quantum time scale, we can neglect the time-dependence of the background metric. Thus we treat the metric components $n$, $a$ and $b$ as functions of only $y$ and the 3D scale factor $a_{\rm sc}$ as a constant, in the rest of the paper.
@@ 3.1.1 Equation of motion
As shown in App. C.1, the Fourier component $\tilde{G}_{\rm B}(p_{\rm 4d}^{2},y,y^{\prime})$ is expressed as
@@
The functions $C_{0,L}(y;p_{\rm 4d}^{2})$ are the basis of solutions to (3.6), and their definitions are provided in Appendix B.1.
@@ 3.1.3 Energy-momentum tensor
$=\frac{a_{\rm sc}^{3}}{4\pi^{3}}\lim_{y^{\prime}\to y}\int_{0}^{\infty}d\tilde{p}_{t}\int_{0}^{\infty}d\tilde{p}_{3}\;\tilde{p}_{3}^{2}\left[-\tilde{p}_{t}^{2}+\tilde{p}_{\rm 3d}^{2}+\frac{n^{2}(y)}{b^{2}(y)}\partial_{y}\partial_{y}^{\prime}+n^{2}(y)M^{2}\right]\tilde{G}_{\rm B}(\tilde{p}_{4}^{2},y,y^{\prime}),$ (3.16)
@@ 3.2.1 Equations of motion
As shown in App. C.2, the components of the 5D propagator have the following forms.
@@
where the functions $\bar{G}_{ab}(p_{\rm 4d}^{2},y,y^{\prime})$ ($a,b=\chi,\lambda$) are read off from (C.16), which are expressed in terms of the basis functions for $\cal O$$_{\pm}$ defined in Appendix B.2.
@@ 4 Cutoff-dependence
Note that the expressions in (3.18), (3.19) and (3.39) diverges by taking the limit of $y^{\prime}\to y$. Since we are interested in the dependence on the cutoff scale $\Lambda_{\rm cut}$, we keep $\Lambda_{\rm cut}\equiv(y^{\prime}-y)^{-1}$ finite. [^8] Since a higher-dimensional field theory is non-renormalizable, it should be regarded as effective theory that is valid only below some cut-off energy scale $\Lambda_{\rm cut}$. We should also notice that there is an ambiguity in a way of introducing $\Lambda_{\rm cut}$ to the expressions of $\left\langle T_{MN}\right\rangle$. For example, (3.18) can be regularized as
@@ 4.1 Scalar sector
The behavior of the 5D propagator at large $\rho$ is shown in Appendix D.1. Using (D.12) and (D.14), we have
@@
Here ${\cal H}_{l}(y,y^{\prime})$ and ${\cal S}_{l}(y,y^{\prime})$ are defined in Appendix D.1, and are determined by the background metric. Therefore, the divergent parts $\left\langle T_{MN}^{\rm b}\right\rangle^{\rm UV}$ are expressed as
@@
and $\Gamma(a,x)\equiv\int_{x}^{\infty}dw\;e^{-w}w^{a-1}$ is the incomplete gamma function. Using the expansion (D.35), we obtain
@@ 4.2 Spinor sector
From this and (D.28), the $\Lambda_{\rm cut}$-dependent part of $\left\langle T^{\rm f}_{tt}\right\rangle$ is
@@
Using the expansion (D.35), we obtain
@@ 4.4 Finite part
where $H_{u\chi}(y;\rho^{2})$ and $H_{u\lambda}(y;\rho^{2})$ ($u=0,L$) are defined in (C.17).
@@ 5 Summary and discussions
The statement 4 indicates that $\left\langle T_{MN}\right\rangle$ cannot contribute to the renormalization of the 5D cosmological constant $\Lambda_{\rm cc}^{\rm(5D)}$ in a general background other than AdS or flat spacetime because the $y$-dependence is different from that of $g_{MN}\Lambda_{\rm cc}^{\rm(5D)}$. In that case, $\left\langle T_{MN}\right\rangle$ have large ${\cal O}(\Lambda_{\rm cut}^{3})$ values. From the viewpoint of the energy cost, it seems natural that the geometry of the extra dimension is forced to be the flat (or AdS) space, depending on the mass relations, through the time evolution governed by (3.3). In other words, the dynamics hides the explicit $\Lambda_{\rm cut}$-dependence of the vacuum energy. If this is the case, the geometry is determined by the mass spectrum, and the quantum contribution to the vacuum energy affects the modulus (radion) stabilization. Thus, it is interesting to check whether such process actually happens. We will discuss this issue by numerically solving the Einstein field equation (3.3) in the subsequent papers.
@@ Appendix A Gamma matrices and covariant derivatives
Appendix B Basis functions
B.1 Scalar sector

### v2

@@ 1 Introduction
The paper is organized as follows. In the next section, we provide a brief review of the calculation of the vacuum energy density and the pressure in 4D effective theory, and derive the conditions required to cancel the $\Lambda_{\rm cut}$-dependent part. In Sec. 3, we calculate the VEV of the 5D energy-momentum tensor $\left\langle T_{MN}\right\rangle$ by using the 5D propagators. In Sec. 4, we discuss the dependence of $\left\langle T_{MN}\right\rangle$ on the cutoff energy scale $\Lambda_{\rm cut}$, and its cancellation among bosons and fermions. Sec. 5 is devoted to the summary and discussions. We have collected our notations in App. A, the explicit component expressions of the Riemann tensor and the Einstein equation for our metric ansatz in App. B, the definitions of the basis functions for solutions to the mode equations in App. C, and the derivation of the 5D propagators in App. D. In App. E, we derive approximate expressions of various quantities for large momentum.
@@ 2 Vacuum energy density in 4D effective theory
When we consider the cosmological evolution of the spacetime with a large extra dimension, such as the dark dimension scenario, we need to solve the 5D field equations because the temperature of the universe can be higher than the KK scale, which is the cutoff scale of the 4D effective theory. The evolution of the spacetime is determined by the Einstein equation, [^6] We adopt the semiclassical approach [3], in which the spacetime is treated classically, while the quantum effects of matter fields are incorporated via $\left\langle T_{MN}\right\rangle$.
${\cal R}_{MN}-\frac{1}{2}g_{MN}{\cal R}+g_{MN}\Lambda_{\rm cc}^{\rm(5D)}$ $=\kappa_{5}\left\langle T_{MN}\right\rangle,$ (3.1)
where ${\cal R}_{MN}$ and ${\cal R}\equiv g^{MN}{\cal R}_{MN}$ are the 5D Ricci tensor and scalar, respectively, $\Lambda_{\rm cc}^{\rm(5D)}$ is the 5D cosmological constant, and $\kappa_{5}$ is the 5D gravitational coupling, and $\left\langle T_{MN}\right\rangle$ is the expectation value of the energy-momentum tensor $T_{MN}$ for quantum fields. In the following, we calculate $\left\langle T_{MN}\right\rangle$ for scalar and spinor fields.
Since the background space expands slowly enough compared to the quantum time scale, we can neglect the time-dependence of the background metric in the calculation of $\left\langle T_{MN}\right\rangle$. Thus, in the following calculations, the 5D metric is parameterized as
$ds^{2}$ $=g_{MN}dx^{M}dx^{N}=-n^{2}(y)dt^{2}+a^{2}(y)\sum_{i=1}^{3}(dx^{i})^{2}+b^{2}(y)dy^{2},$ (3.2)
where $t\equiv x^{0}$, $y\equiv x^{4}$, and $0\leq y\leq L\equiv\pi R$. To simplify the calculations, we further assume that
$a(y)$ $=a_{\rm sc}n(y),$ (3.3)
where $a_{\rm sc}$ is a constant. Although $a_{\rm sc}$ can always be set to 1 by the redefinition of the 3D coordinates $x^{i}$ ($i=1,2,3$), we leave it as an arbitrary positive constant because it becomes the scale factor when we revive the time-dependence of the background metric. In fact, when we solve the evolution equation (3.1), we should revive the time-dependence of $n$, $a_{\rm sc}$ and $b$ in the our results (3.18), (3.19) and (3.39). The explicit forms of the evolution equations are given by (B.5) with (B.6) and (B.7) in Appendix B.2.
@@ 3.1.1 Equation of motion
As shown in App. D.1, the Fourier component $\tilde{G}_{\rm B}(p_{\rm 4d}^{2},y,y^{\prime})$ is expressed as
@@
The functions $C_{0,L}(y;p_{\rm 4d}^{2})$ are the basis of solutions to (3.6), and their definitions are provided in Appendix C.1.
@@ 3.1.3 Energy-momentum tensor
$=\frac{a_{\rm sc}^{3}}{4\pi^{3}}\lim_{y^{\prime}\to y}\int_{0}^{\infty}d\tilde{p}_{t}\int_{0}^{\infty}d\tilde{p}_{\rm 3d}\;\tilde{p}_{\rm 3d}^{2}\left[-\tilde{p}_{t}^{2}+\tilde{p}_{\rm 3d}^{2}+\frac{n^{2}(y)}{b^{2}(y)}\partial_{y}\partial_{y}^{\prime}+n^{2}(y)M^{2}\right]\tilde{G}_{\rm B}(\tilde{p}_{4}^{2},y,y^{\prime}),$ (3.16)
@@ 3.2.1 Equations of motion
As shown in App. D.2, the components of the 5D propagator have the following forms.
@@
where the functions $\bar{G}_{ab}(p_{\rm 4d}^{2},y,y^{\prime})$ ($a,b=\chi,\lambda$) are read off from (D.16), which are expressed in terms of the basis functions for $\cal O$$_{\pm}$ defined in Appendix C.2.
@@ 4 Cutoff-dependence
Note that the expressions in (3.18), (3.19) and (3.39) diverges by taking the limit of $y^{\prime}\to y$. Since we are interested in the dependence on the cutoff scale $\Lambda_{\rm cut}$, we keep $\Lambda_{\rm cut}\equiv(y^{\prime}-y)^{-1}$ finite. [^8] In our setup, we have neglected any interaction terms because they are irrelevant to the one-loop contributions to the vacuum energy density and the pressures. Thus the theory is renormalizable. However, once interactions are turned on, it becomes non-renormalizable and should be regarded as effective theory that is valid only below some cut-off energy scale $\Lambda_{\rm cut}$. We should also notice that there is an ambiguity in a way of introducing $\Lambda_{\rm cut}$ to the expressions of $\left\langle T_{MN}\right\rangle$. For example, (3.18) can be regularized as
@@ 4.1 Scalar sector
The behavior of the 5D propagator at large $\rho$ is shown in Appendix E.1. Using (E.12) and (E.14), we have
@@
Here ${\cal H}_{l}(y,y^{\prime})$ and ${\cal S}_{l}(y,y^{\prime})$ are defined in Appendix E.1, and are determined by the background metric. Therefore, the divergent parts $\left\langle T_{MN}^{\rm b}\right\rangle^{\rm UV}$ are expressed as
@@
and $\Gamma(a,x)\equiv\int_{x}^{\infty}dw\;e^{-w}w^{a-1}$ is the incomplete gamma function. Using the expansion (E.35), we obtain
@@ 4.2 Spinor sector
From this and (E.28), the $\Lambda_{\rm cut}$-dependent part of $\left\langle T^{\rm f}_{tt}\right\rangle$ is
@@
Using the expansion (E.35), we obtain
@@ 4.4 Finite part
where $H_{u\chi}(y;\rho^{2})$ and $H_{u\lambda}(y;\rho^{2})$ ($u=0,L$) are defined in (D.17).
@@ 5 Summary and discussions
As noted in the footnote 8, the theory considered here is renormalizable because interaction terms are neglected in our setup. Hence all the contributions with positive powers in $\Lambda_{\rm cut}$ should be absorbed into some parameters. The contributions to $\left\langle T_{MN}\right\rangle$ that have the same $y$-dependence as $g_{MN}$ can be absorbed to the 5D cosmological constant $\Lambda_{\rm cc}^{\rm(5D)}$. However ,the results in Sec. 4.3 show that the cosmological constant term is not enough to renormalize all the divergent contributions. Therefore, we expect that such contributions are absorbed to some higher order terms in the background curvature terms, such as ${\cal R}^{MNLP}{\cal R}_{MNLP}$ or ${\cal R}^{MN}{\cal R}_{MN}$, which will be induced at loop level. In fact, in the case of $\left\{n(y)=1+cy^{2},\;b(y)=b_{\rm c}\right\}$, $C_{1}^{(tt)}(y)$ and $C_{1}^{(yy)}(y)$ contain terms proportional to $(1+cy^{2})^{-1}$ and $(1+cy^{2})^{-2}$. Such terms can arise from ${\cal R}^{MNLP}{\cal R}_{MNLP}$, ${\cal R}^{MN}{\cal R}_{MN}$ or ${\cal R}^{2}$ as shown in (B.8).
We should also note that no $\Lambda_{\rm cut}^{4}$- and $\Lambda_{\rm cut}^{2}$-contributions would appear in the momentum cutoff regularization, instead of the point-splitting regularization we have used. Therefore, such terms may be understood as artificial ones due to an inappropriate choice of the regularization. There may be a specific choice of introducing $\Lambda_{\rm cut}$ in the expressions of $\left\langle T_{MN}\right\rangle$ for the ambiguity mentioned at the beginning of Sec. 4, with which the $\Lambda_{\rm cut}^{4}$- and $\Lambda_{\rm cut}^{2}$-contributions will disappear.
@@
The statement 4 indicates that $\left\langle T_{MN}\right\rangle$ have large ${\cal O}(\Lambda_{\rm cut}^{3})$ values in a general background other than AdS or flat spacetime. From the viewpoint of the energy cost, it seems natural that the geometry of the extra dimension is forced to be the flat (or AdS) space, depending on the mass relations, through the time evolution governed by (3.1). In other words, the dynamics hides the explicit $\Lambda_{\rm cut}$-dependence of the vacuum energy. If this is the case, the geometry is determined by the mass spectrum, and the quantum contribution to the vacuum energy affects the modulus (radion) stabilization. It is interesting to check whether such process actually happens. We will discuss this issue by numerically solving the Einstein field equation (3.1) in the subsequent papers.
@@ Appendix A Gamma matrices and covariant derivatives
Appendix B Riemann tensor and Einstein equations
B.1 Riemann and Ricci tensors
Under the metric ansatz (3.2), the nonvanishing components of the Riemann tensor are
${\cal R}_{titj}$ $=\left(-a\ddot{a}+\frac{\dot{n}a\dot{a}}{n}+\frac{nn^{\prime}aa^{\prime}}{b^{2}}\right)\delta_{ij},$
${\cal R}_{tijy}$ $=\left(a\dot{a}^{\prime}-\frac{n^{\prime}a\dot{a}}{n}-\frac{aa^{\prime}\dot{b}}{b}\right)\delta_{ij},$
${\cal R}_{tyty}$ $=-b\ddot{b}+\frac{\dot{n}b\dot{b}}{n}+nn^{\prime\prime}-\frac{nn^{\prime}b^{\prime}}{b},$
${\cal R}_{ijkl}$ $=a^{2}\left(\frac{\dot{a}^{2}}{n^{2}}-\frac{a^{\prime 2}}{b^{2}}\right)\left(\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}\right),$
${\cal R}_{yiyj}$ $=\left(-aa^{\prime\prime}+\frac{aa^{\prime}b^{\prime}}{b}+\frac{a\dot{a}b\dot{b}}{n^{2}}\right)\delta_{ij},$ (B.1)
and components related to these by the index symmetries, and the nonvanishing components of the Ricci tensor are
${\cal R}_{tt}$ $=-\frac{3\ddot{a}}{a}-\frac{\ddot{b}}{b}+\left(\frac{3\dot{a}}{a}+\frac{\dot{b}}{b}\right)\frac{\dot{n}}{n}+\frac{nn^{\prime\prime}}{b^{2}}+\left(\frac{3a^{\prime}}{a}-\frac{b^{\prime}}{b}\right)\frac{nn^{\prime}}{b^{2}},$
${\cal R}_{ty}$ $=-\frac{3\dot{a}^{\prime}}{a}+\frac{3n^{\prime}\dot{a}}{na}+\frac{3a^{\prime}\dot{b}}{ab},$
${\cal R}_{ij}$ $=\left[\frac{a\ddot{a}}{n^{2}}-\frac{aa^{\prime\prime}}{b^{2}}-\left(\frac{\dot{n}}{n}-\frac{2\dot{a}}{a}-\frac{\dot{b}}{b}\right)\frac{a\dot{a}}{n^{2}}-\left(\frac{n^{\prime}}{n}+\frac{2a^{\prime}}{a}-\frac{b^{\prime}}{b}\right)\frac{aa^{\prime}}{b^{2}}\right]\delta_{ij},$
${\cal R}_{yy}$ $=\frac{b\ddot{b}}{n^{2}}-\frac{n^{\prime\prime}}{n}-\frac{3a^{\prime\prime}}{a}-\left(\frac{\dot{n}}{n}-\frac{3\dot{a}}{a}\right)\frac{b\dot{b}}{n^{2}}+\left(\frac{n^{\prime}}{n}+\frac{3a^{\prime}}{a}\right)\frac{b^{\prime}}{b}.$ (B.2)
B.2 Einstein equations
Utilizing the expressions in (B.2), the Einstein equation (3.1) is summarized as
$\frac{\ddot{a}}{a}$ $=\frac{\dot{a}}{a}\left(\frac{\dot{n}}{n}-\frac{\dot{a}}{a}\right)+\frac{n^{2}}{b^{2}}\frac{a^{\prime}}{a}\left(\frac{n^{\prime}}{n}+\frac{a^{\prime}}{a}\right)-\frac{\kappa_{5}n^{2}}{3b^{2}}\left\langle T_{yy}\right\rangle+\frac{n^{2}}{3}\Lambda_{5},$
$\frac{\ddot{b}}{b}$ $=\frac{4\dot{a}^{2}}{a^{2}}+\frac{\dot{b}}{b}\left(\frac{\dot{n}}{n}+\frac{\dot{a}}{a}\right)+\frac{n^{2}}{b^{2}}\left[\frac{n^{\prime\prime}}{n}-\frac{a^{\prime\prime}}{a}-\frac{4a^{\prime 2}}{a^{2}}-\frac{b^{\prime}}{b}\left(\frac{n^{\prime}}{n}-\frac{a^{\prime}}{a}\right)\right]$
$\quad-\kappa_{5}\left(\left\langle T_{tt}\right\rangle+\frac{n^{2}}{a^{2}}\left\langle T_{\rm 3d}\right\rangle-\frac{2n^{2}}{3b^{2}}\left\langle T_{yy}\right\rangle\right)-\frac{2n^{2}}{3}\Lambda_{5},$ (B.3)
where $\left\langle T_{ij}\right\rangle\equiv\delta_{ij}\left\langle T_{\rm 3d}\right\rangle$, with
$\frac{\dot{a}}{a}\left(\frac{\dot{a}}{a}+\frac{\dot{b}}{b}\right)-\frac{n^{2}}{b^{2}}\left[\frac{a^{\prime\prime}}{a}+\frac{a^{\prime}}{a}\left(\frac{a^{\prime}}{a}-\frac{b^{\prime}}{b}\right)\right]$ $=\frac{\kappa_{5}}{3}\left\langle T_{tt}\right\rangle+\frac{n^{2}}{3}\Lambda_{5},$
$\frac{\dot{a}^{\prime}}{a}$ $=\frac{n^{\prime}\dot{a}}{na}+\frac{a^{\prime}\dot{b}}{ab}.$ (B.4)
The equations in (B.4) do not contain the second time derivatives, and thus should be understood as constraints on the initial conditions. [^10] Once the equations in (B.4) are satisfied at the initial time, they remain to be hold at later times.
If we restrict the metric components as (3.3), the above equations become
$\frac{\ddot{n}}{n}$ $=-\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\left(\frac{3\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\right)-\frac{\ddot{a}_{\rm sc}}{a_{\rm sc}}+\frac{2n^{\prime 2}}{b^{2}}-\frac{\kappa_{5}n^{2}}{3b^{2}}\left\langle T_{yy}\right\rangle+\frac{n^{2}}{3}\Lambda_{5},$
$\frac{\ddot{b}}{b}$ $=4\left(\frac{\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\right)^{2}+\frac{\dot{b}}{b}\left(\frac{2\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\right)-\frac{4n^{\prime 2}}{b^{2}}+\frac{2\kappa_{5}n^{2}}{3b^{2}}\left\langle T_{yy}\right\rangle-\frac{2n^{2}}{3}\Lambda_{5},$ (B.5)
with
$\left(\frac{\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\right)\left(\frac{\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}+\frac{\dot{b}}{b}\right)-\frac{n^{2}}{b^{2}}\left[\frac{n^{\prime\prime}}{n}+\frac{n^{\prime}}{n}\left(\frac{n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)\right]$ $=\frac{\kappa_{5}}{3}\left\langle T_{tt}\right\rangle+\frac{n^{2}}{3}\Lambda_{5},$ (B.6)
and
$\frac{\dot{n}^{\prime}}{n}$ $=\frac{n^{\prime}}{n}-\frac{n^{\prime}\dot{a}_{\rm sc}}{na_{\rm sc}}+\left(\frac{\dot{n}}{n}+\frac{\dot{a}_{\rm sc}}{a_{\rm sc}}\right)^{-1}\frac{n^{\prime}\dot{b}}{nb}.$ (B.7)
B.3 Higher curvature terms
Under the assumption (3.3) and neglecting the $t$-dependence of the background, we have
${\cal R}^{MNLP}{\cal R}_{MNLP}$ $=\frac{8}{b^{4}}\left[\frac{3n^{\prime 4}}{n^{4}}+2\left(\frac{n^{\prime\prime}}{n}-\frac{n^{\prime}b^{\prime}}{nb}\right)^{2}\right],$
${\cal R}^{MN}{\cal R}_{MN}$ $=\frac{4}{b^{4}}\left[5\left(\frac{n^{\prime\prime}}{n}-\frac{n^{\prime}b^{\prime}}{nb}\right)^{2}+\frac{6n^{\prime 2}}{n^{2}}\left(\frac{n^{\prime\prime}}{n}-\frac{n^{\prime}b^{\prime}}{nb}\right)+\frac{9n^{\prime 4}}{n^{4}}\right],$
${\cal R}^{2}$ $=\frac{64}{b^{4}}\left(\frac{n^{\prime\prime}}{n}-\frac{n^{\prime}b^{\prime}}{nb}+\frac{3n^{\prime 2}}{n^{2}}\right)^{2}.$ (B.8)
Appendix C Basis functions
C.1 Scalar sector